OpenStudy (anonymous):

Find the derivative of the given function at the indicated point. f(x)=1/x,a = 2 f′(a)= lim (f(a+h) - f(a))/h h --> 0

OpenStudy (anonymous):

f′(a)= lim (1/(2+h) - .5)/h h --> 0 Im at this part, where do I go from here?

OpenStudy (anonymous):

\[f'(x) =\lim_{h \rightarrow 0} -1/x^2\]

OpenStudy (anonymous):

But how did you get from what I have to what you have?

OpenStudy (anonymous):

\[f'(x)= \lim_{h \rightarrow 0}\frac{ \frac{ 1 }{ x-h }-\frac{ 1 }{ x }}{ h}\]